# Multilingual != Multicultural: Evaluating Gaps Between Multilingual Capabilities and Cultural Alignment in LLMs

### v1

@@ 3.1 Models
This study examines two model families with distinct development approaches: the Gemma family of models (Gemma et al., 2024) and OpenAI’s commercial releases in the turbo-range (OpenAI, 2022; OpenAI et al., 2024a, b). These two families allow us to investigate two different modes of model improvement: scaling (Gemma) and iterative product development with scaling and post-training improvements (OpenAI). Other preliminary experiments included different versions of LLaMA models (Touvron et al., 2023) and Mistral models (Jiang et al., 2023). However, these models either failed to consistently follow instructions or always answered in English regardless of the prompt language.
Gemma:
The Gemma family comprises open-weight models ranging from 2 to 27 billion parameters (Gemma et al., 2024). The smaller models (2B and 9B) are trained through knowledge distillation from a more capable teacher model, while the 27B model uses standard pre-training objectives. While instruction tuning uses primarily English data, evaluations show strong performance on multilingual tasks (Nielsen, 2023). This restricted training regime, combined with systematic variation in model scale, provides a controlled environment for investigating whether cultural alignment capabilities emerge naturally with increased model size (Kaplan et al., 2020).
OpenAI:
OpenAI’s ‘turbo’ series represents iterative commercial development driven by customer requirements. The original gpt-3.5-turbo, based on the InstructGPT architecture (Ouyang et al., 2022), established the foundation for conversational AI. Each subsequent release addresses specific market needs: gpt-4-turbo offers reduced computation costs of the more powerful gpt-4 model (OpenAI et al., 2024a) through distillation, while gpt-4o expands multilingual capabilities to serve a global user base OpenAI et al. (2024b). Part of the purported multilingual improvements of gpt-4o comes from a more representative tokenizer, though this has been critiqued Yang et al. (2024). This commercially-motivated progression, with each model optimized for different operational priorities, provides a complementary perspective to Gemma’s controlled scaling study.
@@ 3.2 RQ1: Multilingual Cultural Alignment
[…]
To statistically assess whether improving the multilingual capabilities of LLMs improves cultural alignment, we construct a linear mixed effects regression model (Luke, 2017) based on the experimental setup described above. Here, we relate the cultural alignment of the model to the improvements in multilingual capabilities controlled by relevant factors:
$\begin{split}\text{CA}_{i}&\sim\mathcal{N}(\mu_{i},\sigma^{2})\\ \mu_{i}&=\alpha_{j[i]}+\beta_{1}X_{cons,i}+\sum_{f,l}\beta_{flm}X_{fl,i}X_{m,i}\\ \alpha_{j}&\sim\mathcal{N}(\mu_{\alpha},\sigma^{2}_{\alpha}),\quad\text{for }j=1,\ldots,J\end{split}$ (1)
Here, Eq. 1 models the cultural alignment ($\text{CA}_{i}$) of an LLM, $i$, as a linear mixed-effects model. The mean $\mu_{i}$ includes a random intercept $\alpha_{j}$ for each model, the fixed effect of consistency ($\beta_{1}X_{cons,i}$), and interaction effects ($\sum_{f,l}\beta_{flm}X_{fl,i}X_{m}$) between model family (either ‘OpenAI’ or ‘Gemma’), language (e.g., Portuguese or Danish), and multilingual capability (average score on selected benchmarks; see Appendix F). Here, consistency is an LLM’s stability in value stances across repeated questions, calculated as the Pearson correlation between value polarity scores (defined in §2) of repeated responses to identical topics (see Appendix A). A score of 1.0 indicates perfect consistency, 0.0 indicates randomness. Population-level resampling of human survey responses yields values between 0.66 and 0.84 (see Fig 3). Assumption checks for the regression can be seen in Appendix D.
The above statistical model allows us to analyse the relationship between multilingual capabilities and cultural alignment in model families at the level of individual languages. For instance, we might find an improvement for Gemma models for Danish but not for Dutch or vice versa.
@@ 3.3 RQ2: US-Centric Bias
We analyze model bias by comparing cultural alignment between US and local values, where “local” refers to values in country or countries where a given language is natively spoken. We define US-centric bias as an LLM showing higher cultural alignment with US value distributions compared to local ones. To quantify this bias, we use a linear regression model that measures the differential effect of US versus local value alignment:
@@
The regression’s intercept (i.e., the base case) is a baseline that produces uniformly random value polarity scores. $\mathcal{M}$ is the set of models and $\mathcal{L}$ is the set of languages. US is a boolean feature denoting whether the cultural alignment is to the US (if $1$) or the local values (if 00). We primarily analyse the coefficients with US ($\beta_{ml}^{US}$) since these provide the partial effect of US-centric bias, i.e., how much more/less a given LLM is aligned to US rather than local values. Assumption checks for the regression can be seen in Appendix D.
@@ 4 Results
4.1 Multilingual Value Alignment (RQ1)
To understand how improving multicultural capabilities affects cultural alignment, we must first examine the stability of LLMs’ cultural values. When evaluating LLMs that lack stable internal values, apparent improvements in cultural alignment may simply reflect reduced response variance rather than genuine advances in cultural alignment (Röttger et al., 2024; Kahneman et al., 2021). We therefore analyse both the self-consistency of LLM responses and how alignment changes with model improvements.
@@ 4.1 Multilingual Value Alignment (RQ1)
We find consistently low self-consistency scores across all models and languages compared to human responses in the WVS data (Fig. 3). For WVS data, self-consistency scores (measured as Pearson correlations between bootstrapped samples of responses) range from 0.66 to 0.84 across countries (see Fig 3). In contrast, LLM responses to the same questions under different prompt variations show significantly lower consistency, even in English where performance is highest due to English-dominated training data (OpenAI et al., 2024a; Gemma et al., 2024).
This lower self-consistency complicates our cultural alignment analysis Wright et al. (2024). Drawing on Kahneman et al. (2021)’s noise framework, we recognise that inconsistent responses can be as detrimental as bias. To address the noise, we employ larger sample sizes and incorporate consistency measures in our regression analyses.
@@
The relationship between model improvements and cultural alignment varies substantially across languages and model families (Fig. 1). For Gemma, there is a significant positive relationship between multilingual capabilities and cultural alignment for all languages. In contrast, OpenAI only has a significant and positive relationship for English ($\beta_{\text{OpenAI},\text{en}}=0.11,p=0.04$), whereas for Danish and Portuguese the relationship is insignificant ($\beta_{\text{OpenAI},\text{pt}}=0.089,p=0.10$; $\beta_{\text{OpenAI},\text{da}}=0.009,p=0.89$) and for Dutch the effect is significant and negative ($\beta_{\text{OpenAI},\text{nl}}=-0.19,p=0.004$).
A capability threshold may explain the discrepancy between multilingual performance and cultural alignment. Up to a point, multilingual capabilities might improve alignment by enhancing cultural knowledge and instruction-following. Beyond this threshold, other factors likely become more influential (Kirk et al., 2024a). This could explain the stronger relationship in Gemma versus OpenAI’s LLMs (see Fig 4). Future work with the Qwen-2.5 family (Qwen et al., 2025), which range from 500M to 72B parameters, could help validate this hypothesis.
Looking at specific languages, we find nuanced patterns of improvement in cultural representations across model versions. Portuguese shows consistent improvements across both model families, particularly for local (Portuguese and Brazilian) values compared to global values. Danish exhibits positive but non-significant trends, while Dutch shows more variable results—including a notable improvement only with GPT-4o for OpenAI models (See fig. 6). For a full breakdown of language performance see Appendix C.
Furthermore, the dominant effect of self-consistency ($\beta_{\text{consistency}}=0.62,p\ll 0.001$) compared to multilingual capability suggests that noise remains a major limiting factor in analysing cultural alignment. This aligns with broader findings about the instability of LLM value elicitation (Röttger et al., 2024). Moreover, even the highest observed alignment scores (around 0.4; see Fig 4) indicate substantial room for improvement in how well LLMs match human cultural values and behaviours.
In conclusion, our analysis reveals a complex relationship between model improvements and cultural alignment. While some languages show progressive improvements in cultural alignment from model scaling or iterative commercial development, others show minimal or inconsistent improvements. These findings, combined with the generally low self-consistency of LLM responses, suggest that improving multilingual model capabilities does not automatically lead to better cultural alignment across all languages and contexts.
@@ 4.2 US-centric Bias (RQ2)
Figure 5: US-centric bias coefficients across LLMs and languages (see Eq. 2). Error bars are standard errors from the regression. Positive values indicate bias.
Our analysis reveals distinct patterns of US-centric bias across both languages and model families (Fig. 5). Languages show different susceptibilities to US bias: five out of six LLMs exhibit US-centric bias in Dutch, all in English, none in Portuguese, and Danish falls in between with three of the six LLMs showing bias.
Within model families, we observe varying patterns of changes in US-centric bias. The Gemma models show a clear trajectory in Dutch, becoming significantly less biased in successive progressions. However, this improvement isn’t universal: for Portuguese and English, there are no significant differences in bias between progressions within either family. For Danish, early models (gpt-3.5-turbo and gemma-2-2b-it) exhibit US-centric bias, while later progressions show reduced bias, particularly in the OpenAI family.
In conclusion, we find significant US-centric bias in certain LLM-language combinations, particularly in Dutch and Danish. However, the progression of this bias across model versions varies by language and model family. While some improvements are visible (e.g., Gemma models’ reduced bias in Dutch), the overall pattern suggests that advancing model capabilities does not consistently reduce or increase US-centric bias.
Figure 6: Cultural alignment scores for LLM responses in Danish (top) and Dutch (bottom). Left: comparison with English speakers; middle: Denmark/Netherlands vs US; right: alignment with global values. Error bars show bootstrapped CIs.
@@ 5 Related Work
Recent work emphasizes the need for systematic auditing of LLMs’ cultural alignment, particularly as these models are deployed globally (Kirk et al., 2024a; Mökander et al., 2024; Kirk et al., 2024b). Prior empirical approaches have primarily taken two paths: using tranformations based on Hofstede’s cultural dimensions framework or directly comparing against survey responses. Studies using Hofstede’s dimensions (Masoud et al., 2025; Cao et al., 2023) provide structured cross-cultural comparisons through latent variable analysis. However, these studies assume that LLMs’ latent dimensions map directly onto human dimensions, since they use formulas calibrated for humans—an assumption that warrants scrutiny (Shanahan, 2024).
Recent work has explored using LLMs to simulate survey responses for assessing cultural alignment (Tao et al., 2024; AlKhamissi et al., 2024). Prior approaches focused on individual-level responses. In contrast, our method generates distributions of opinions across hypothetical survey participants, enabling direct comparison with population-level statistics. This distribution-based approach offers three key advantages. First, it better captures the inherent variation in cultural values within populations, paving the way for investigating distributional alignment (Sorensen et al., 2024). Second, it enables principled statistical comparison against large-scale survey data like the World Values Survey (EVS/WVS, 2022). Finally, the framework is easy to extend to new languages by automatically translating the prompts. We detail our quantitative framework for measuring alignment with observed population distributions in §2.
There is also an increasing body of work investigating political biases in LLMs Röttger et al. (2024, 2025); Hartmann et al. (2023). Much of this work also relies on human political surveys like the Political Compass Test. However, recent work has called for increased attention to the randomness inherent in LLM decoding at non-zero temperatures can create instability in attributes Röttger et al. (2024); Wright et al. (2024). We expand on this work by including multilingual perspectives, and constructing prompts on a wide range of prompt variations (see §2). These prompt variations combined with statistically accounting for self-consistency in our statistical analysis (see §3.2) allow us to get a more robust measure of cultural alignment.
@@
Previous work has focused primarily on English-language performance (Tao et al., 2024) or a single LLM (OpenAI et al., 2024a; Cao et al., 2023). Our work extends this by examining how cultural alignment varies across both model families and languages, providing insight into how different development approaches—scaling and commercial product development—influence cultural representation capabilities.
There is already positive progress on improving the cross-cultural participation in alignment. Two notable projects are PRISM and AYA (Kirk et al., 2024b; Üstün et al., 2024). PRISM is a large dataset of conversational preferences from a diverse participant pool. While the data is predominantly in English, it could be an important resource for better understanding and modelling diverse cultural preferences. The AYA dataset is a massively multilingual instruction fine-tuning dataset. Through improvement non-English instruction tuning, AYA could prove a central component in decreasing noise in LLM responses facilitating cultural alignment.
@@ 6 Conclusion
Increased multilingual capabilities do not guarantee improved cultural alignment in Large Language Models. Through systematic comparison of two model families—Gemma and OpenAI—we find that the relationship between improvements in multilingual capability and cultural alignment is complex. While some languages show clear improvements in alignment with increased model capabilities (e.g., Portuguese), others exhibit inconsistent patterns, suggesting that cultural alignment does not automatically follow gains in multilingual capabilities. Our distribution-matching methodology using World Values Survey data enabled the detection of these nuanced patterns across languages and cultural contexts.

### v2

@@ 3.1 Models
We examine three model families representing different development approaches: Gemma Gemma et al. (2024) and OLMo (OLMo et al., 2025) for improvements through scaling and OpenAI’s turbo series for commercial product development, combining scaling with post-training improvements (OpenAI, 2022; OpenAI et al., 2024a, b). Other preliminary experiments included different versions of LLaMA models (Touvron et al., 2023) and Mistral models (Jiang et al., 2023). However, these models either failed to consistently follow instructions or always answered in English regardless of the prompt language. See the supplementary materials for a more thorough description of the LLMs.
@@ 3.2 RQ1: Multilingual Cultural Alignment
Figure 4: Language capability (x-axis) vs cultural alignment scores (y-axis) across languages. Stars indicate significance ($p<.05$) in our linear mixed-effects regression of multiple runs (See §3.2). OpenAI models (blue) and OLMo models (red) show negative/insignificant relationships outside of English, while the Gemma models (green) show positive relationships throughout ($p<.05$).
To statistically assess whether improving the multilingual capabilities of LLMs improves cultural alignment, we construct a linear mixed-effects regression (LMER; Luke, 2017) based on the experimental setup described above. Our LMER follows standard practices and has three core components:
  • Core coefficient: The coefficient of interest is the three-way interaction between model family, language, and multilingual capability. This tests whether the multilingual capability–alignment relationship differs by model family and response language, directly addressing RQ1.
  • Random effects: We include a model-specific random intercept $\alpha_{j}$ to account for repeated measures of cultural alignment for the same LLM. This models variation between LLMs and can improve efficiency over standard linear regressions (Luke, 2017).
  • Control for self-consistency: We include a consistency-by-language term to help ensure that higher alignment scores reflect genuine cultural adaptation rather than reduced response noise, which can inflate scores (Kahneman et al., 2021).
We calculate self-consistency as the Spearman correlation between value polarity scores (defined in §2) of repeated responses to identical topics, adjusted by the reliability of the LLM annotation (see §2.2; Charles, 2005). A score of 1.0 indicates perfect consistency; 0.0 indicates random responses. Population-level resampling of the human WVS responses yields values between 0.66 and 0.84 (see Fig.3 and the supplementary materials).
Formally, the model is specified in Eq. 3:
$\begin{split}\text{CA}_{i}&\sim\mathcal{N}(\mu_{i},\sigma^{2}),\\ \mu_{i}&=\alpha_{j[i]}+\beta_{1l}\,X_{\text{cons},i}X_{l,i}\\ &\quad+\beta_{flm}\,X_{m,i}X_{f,i}X_{l,i},\\ \alpha_{j}&\sim\mathcal{N}(\mu_{\alpha},\sigma_{\alpha}^{2}),\quad j=1,\ldots,J.\end{split}$ (3)
where $i$ indexes responses and $j[i]$ denotes the LLM producing response $i$. Here $X_{\text{cons},i}$ is the self-consistency score for response $i$, $X_{l,i}$ is the set of language indicators, $X_{f,i}$ is the set of model-family indicators, and $X_{m,i}$ is the multilingual capability score. The residual variance $\sigma^{2}$ represents within-LLM variation in alignment scores not explained by the fixed effects or model-specific intercept, while $\sigma_{\alpha}^{2}$ represents between-LLM variation in average alignment.
The above statistical model allows us to analyse the relationship between multilingual capabilities and cultural alignment in model families at the level of individual languages. For example, we might find that multilingual capabilities improve cultural alignment for Gemma models for Danish but not for Dutch or vice versa.
@@ 3.3 RQ2: US-Centric Bias
We analyse model bias by comparing cultural alignment between US and local values, where “local” refers to values in the country or countries where a given language is natively spoken. We define US-centric bias as an LLM showing higher cultural alignment with US value distributions compared to local ones. To quantify this bias, we use a linear regression model that measures the differential effect of US versus local value alignment:
@@
The regression’s intercept ($\beta_{0}$,i.e., the base case) is a baseline that produces uniformly random value polarity scores. $\mathcal{M}$ is the set of models and $\mathcal{L}$ is the set of languages. US is a boolean feature denoting whether the cultural alignment is to the US (if $1$) or the local values (if 0). We primarily analyse the coefficients with US ($\beta_{ml}^{US}$) since these provide the partial effect of US-centric bias, i.e., how much more/less a given LLM is aligned to US rather than local values. Assumption checks for the regression can be seen in the supplementary materials.
@@ 4 Results
4.1 Multilingual Cultural Alignment (RQ1)
We first examine the stability of LLMs’ cultural values. For LLMs lacking stable internal values, apparent improvements in cultural alignment may reflect reduced response variance rather than genuine advances (Röttger et al., 2024; Kahneman et al., 2021). We therefore analyse both the self-consistency of LLM responses and how alignment changes with model improvements.
@@ 4.1 Multilingual Cultural Alignment (RQ1)
We find low self-consistency scores across all models and languages compared to human responses in the WVS data (Fig. 3). In contrast, LLMs show generally lower self-consistency compared to the human responses, even in English, where instruction-following capabilities are strongest due to English-dominated training data. (OpenAI et al., 2024a; Gemma et al., 2024; OLMo et al., 2025).
This lower self-consistency complicates our cultural alignment analysis Wright et al. (2024). Drawing on Kahneman et al. (2021)’s noise framework, we recognise that inconsistent responses can be as detrimental as bias with respect to the accuracy of the analysis. To address the noise, we employ larger sample sizes and incorporate consistency controls in our regression analyses.
@@
The relationship between model improvements and cultural alignment varies substantially across languages and model families (Fig. 1). For Gemma, there is a strong and significant positive relationship between multilingual capabilities and cultural alignment for all languages. In contrast, the relationships for the GPT-Turbo models are either insignificant or negative. For Dutch and Danish the relationships are insignificant ($\beta_{\text{gpt},\text{nl}}=0.049,p=0.589$$\beta_{\text{gpt},\text{da}}=0.053,p=0.522$), and for Portuguese and English the effect is significant and negative ($\beta_{\text{gpt},\text{en}}=-0.24,p=0.009$, $\beta_{\text{gpt},\text{pt}}=-0.30,p<0.001$). Similarly for OLMo, the relationship is positive for Danish and Dutch ($\beta_{\text{OLMo},\text{da}}=0.44,p<0.001$, $\beta_{\text{OLMo},\text{nl}}=0.29,p<0.001$) and insignificant for English and Portuguese ($\beta_{\text{OLMo},\text{en}}=0.068,p=0.115$, $\beta_{\text{OLMo},\text{pt}}=0.008,p=0.825$).
The mismatch between multilingual performance and cultural alignment could suggest a capability threshold: multilingual improvements might provide rudimentary instruction following skills Nie et al. (2024), but beyond a point, other factors—such as the preferences of developers and annotators—dominate (Kirk et al., 2024b). This could explain the smaller open weights models’ higher coefficients than the gpt-turbo models (see Fig. 4 or Fig. 1). Further work is needed to understand alignment at the sub-national level.
Furthermore, the strong effect of self-consistency ($0.405<\beta_{\text{consistency}}<0.723,p\ll 0.001$) compared to multilingual capability suggests that noise remains a major limiting factor in analysing cultural alignment. This aligns with broader findings about the instability of LLM value elicitation (Röttger et al., 2024; Khan et al., 2025). Moreover, even the highest observed alignment scores (around 0.7; see Fig 4) indicate substantial room for improvement in how well LLMs match human cultural values and behaviours.
In conclusion, our analysis reveals a complex relationship between model improvements and cultural alignment. Although some languages show progressive improvements in cultural alignment from model scaling or iterative commercial development, others show minimal or inconsistent improvements. These findings, combined with the relatively low self-consistency of LLM responses, demonstrate that improved multilingual capability does not guarantee better cultural alignment.
@@ 4.2 US-centric Bias (RQ2)
Figure 5: US-centric bias coefficients across LLMs and languages ($\beta_{BiasUS}$); see Eq. 4). Error bars are standard errors from the regression. Positive values indicate the presence of US-centric bias.
Our analysis reveals distinct patterns of US-centric bias across both languages and model families (Fig. 5). Languages show different susceptibilities to US bias: only one of nine LLMs exhibits US-centric bias in Danish, all in English, all in Portuguese, and none in Dutch. Note that for English, these results mean that the LLM, on average, is relatively more aligned to US values compared to other English-speaking countries like Kenya or the United Kingdom. See the supplementary materials for detailed results.
The overarching pattern is that languages spoken across countries (English and Portuguese) show US-centric bias, whereas languages spoken in only one country (Danish and Dutch) show less US-centric bias. This supports the hypothesis that homogeneity in the training data can counteract US-centric bias—at least for medium-resourced, Western-European languages.
For LLMs, some specific LLMs seem more prone to bias across languages. Specifically, the small gemma-2-2b-it exhibits higher US-centric bias across every language except Dutch. Beyond that, we see no clear progressions in US-centric bias within any family.
In conclusion, language seems a stronger indicator of US-centric bias in LLMs compared to LLM development. Monocultural languages show insignificant to negative bias, while English and Portuguese show significant US-centric bias. Within each LLM family, we find no consistent nor significant change in US-centric bias across LLM versions. These findings underscore the complex relationship between multilingual capability and alignment.
@@ 5 Related Work
Recent work emphasizes the need for systematic auditing of LLMs’ cultural alignment, particularly as these models are deployed globally (Kirk et al., 2024a; Mökander et al., 2024; Kirk et al., 2024b). Prior empirical approaches have primarily taken two paths: using transformations based on Hofstede’s cultural dimensions framework or directly comparing against survey responses. Studies using Hofstede’s dimensions (Masoud et al., 2025; Cao et al., 2023) provide structured cross-cultural comparisons through latent variable analysis. However, these studies assume that LLMs’ latent dimensions map directly onto human dimensions, since they use formulas calibrated for humans—an assumption that warrants scrutiny (Shanahan, 2024; Schröder et al., 2025).
Recent work has explored using LLMs to simulate responses for assessing cultural alignment (Tao et al., 2024; AlKhamissi et al., 2024; Havaldar et al., 2023). Similarly to our work, these works show that LLMs struggle to represent underrepresented personas (AlKhamissi et al., 2024) and emotions (Havaldar et al., 2023) for non-English languages. Prior approaches focused on individual-level responses. In contrast, our method generates distributions of opinions across hypothetical survey participants, enabling direct comparison with population-level statistics. This distribution-based approach offers three key advantages. First, it better captures the inherent variation in cultural values within populations, paving the way for investigating distributional alignment (Sorensen et al., 2024). Second, it enables principled statistical comparison against large-scale survey data like the World Values Survey (EVS/WVS, 2022). Finally, the framework is easy to extend to new languages by automatically translating the prompts. We detail our quantitative framework for measuring alignment with observed population distributions in §2.
There is also an increasing body of work investigating political biases in LLMs Röttger et al. (2024, 2025); Rozado (2024). Much of this work also relies on human political surveys like the Political Compass Test. However, recent work has called for increased attention to how the randomness inherent in LLM decoding at non-zero temperatures can create instability in attributes Röttger et al. (2024); Wright et al. (2024); Khan et al. (2025). We expand on this work by including multilingual perspectives and constructing prompts with a wide range of variations (see §2). These prompt variations, combined with statistically accounting for self-consistency in our statistical analysis (see §3.2), allow us to get a more robust measure of cultural alignment.
@@
Previous work has focused primarily on English-language performance (Tao et al., 2024) or individual LLMs (Arora et al., 2023; Cao et al., 2023). Our work extends this by examining how cultural alignment systematically varies within model families and across languages, providing insight into how different development approaches—scaling and commercial product development—influence cultural representation capabilities.
There is already progress on improving the cross-cultural participation in alignment data. Two notable projects are PRISM and AYA (Kirk et al., 2024b; Üstün et al., 2024). PRISM is a large dataset of conversational preferences from a diverse participant pool. While the data is predominantly in English, it could be an important resource for better understanding and modelling diverse cultural preferences. The AYA dataset is a massively multilingual instruction fine-tuning dataset. AYA could provide further means of realising the demonstrated benefits of multilingual training Nie et al. (2024).
@@ 6 Conclusion
Increased multilingual capabilities do not guarantee improved cultural alignment in Large Language Models. Through systematic comparison of three model families—Gemma, OLMo, and OpenAI’s GPTs—we find that the relationship between improvements in multilingual capability and cultural alignment is complex. While some languages show clear improvements in alignment with increased model capabilities (e.g., Danish), others exhibit inconsistent patterns, suggesting that cultural alignment does not automatically follow gains in multilingual capabilities. Our distribution-matching methodology using World Values Survey data enabled the detection of these nuanced patterns across languages and cultural contexts.
We also find that, contrary to popular discourse, LLMs do not exhibit US-centric bias across all languages; in Danish and Dutch, they align more closely with the values of Denmark and the Netherlands, respectively, than with the US. This fits with the hypothesis that more culturally uniform data leads to less US-centric bias. Both English and Portuguese are spoken in multiple countries, whereas Dutch and Danish are predominantly spoken in one. To further validate this claim, future work could include other multi-cultural languages (like Spanish or Swahili) and monocultural languages (like Japanese)—especially with a wider geographical reach to preclude European bias.